OpenStudy (anonymous):

Graphing Radical Functions (just find domain and range.) Hi! I am working on the study of graphing radical functions. I have already graphed these functions, I just do not understand how to write domain&range for both. I have to write it in terms of "all real numbers" I will include all 5 graphs below. Would really appreciate help.

OpenStudy (anonymous):

Graph 1

OpenStudy (anonymous):

Graph 2

OpenStudy (anonymous):

Graph 3

OpenStudy (anonymous):

Graph 4

OpenStudy (anonymous):

Graph 5

OpenStudy (anonymous):

okay well basically you have to find the asymptote of each function and state that that's the only point that the function does not equal (e.g. all real numbers except"whatever x-intercept the asymptote is at")

OpenStudy (anonymous):

do you know how to find the vertical asymptotes ?

OpenStudy (anonymous):

I am not sure how to do that. I was sadly not introduced throughly to the unit :(

OpenStudy (anonymous):

http://www.purplemath.com/modules/asymtote.htm

OpenStudy (anonymous):

@Hero @pooja195 @triciaal

OpenStudy (anonymous):

Yeah I am not totally understanding how to relate that to finding domain and range i'm a bit confused

OpenStudy (loser66):

For the first one, \(f(x) = \dfrac{1}{x+1}\) take denominator =0, solve for x, what do you get?

OpenStudy (loser66):

knock knock!!! x +1 =0, x =??

OpenStudy (anonymous):

ok! one moment

OpenStudy (anonymous):

-1 ?

OpenStudy (loser66):

yes, domain relate to x, right? but you can't take x =-1 since it makes the function undefined, right? Hence D = \(\mathbb R/{-1}\)

OpenStudy (anonymous):

So that would make the domain for the first graph undefined then?

OpenStudy (loser66):

How about the range? Look at the graph, y goes from -infinitive to +infinitive but at 0, Hence Range is \(\mathbb R/0\)

OpenStudy (loser66):

A little bit difference. domain is all numbers of x that make the function defined. -1 makes the function undefined, hence domain is all real number BUT -1

OpenStudy (anonymous):

ahhhhhhhhhhhhh. That makes sense. I understand. I never understood the "all real numbers except.. ext."

OpenStudy (anonymous):

i'm sorry, I do not understand the notation- by "R/0" you mean that the range is all real numbers but 0?

OpenStudy (anonymous):

I only ask because I have to format it in sentence form

OpenStudy (loser66):

yup

OpenStudy (loser66):

it is not R/0. It is \(\mathbb R/\{0\}\). That the correct notation.

OpenStudy (anonymous):

Yes. My computer would not let me copy that. That is what I meant

OpenStudy (anonymous):

Would you mind helping me with the others?

OpenStudy (loser66):

what is the second function? \(f(x) = \dfrac{1}{x-2}+1\)??

OpenStudy (anonymous):

x would equal 2 here. would this mean that the domain is all real numbers except for 2?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

and how would I find the range? i did not totally understand how you did that for the first graph

OpenStudy (loser66):

to find the range of a radical function, you need asymptotes, that are the line the graph is never touch. like the graph of \(f(x) =\dfrac{1}{x-2}+1\) it looks like